OpenStudy (anonymous):

use linear approximations to estimate the quantity. 4root 16.04

OpenStudy (anonymous):

\[\sqrt[4]{16.04}\]

OpenStudy (zarkon):

let \[f(x)=x^{1/4}\] then \[L(x)=f(16)+f'(16)(x-16)\]

OpenStudy (unklerhaukus):

~\[\sqrt[4]{16}+\sqrt[4]{0.04}\] \[=2+\sqrt[4]{{1}\over{25}}\] \[=2+\sqrt{1\over5}\] \[=2+ {√5 \over 5}\] \[={10+√5 \over 5}\] for some reason what ever i have done is terribly inaccurate

OpenStudy (zarkon):

\[f'(x)=\frac{1}{4}x^{-3/4}\] \[f'(16)=\frac{1}{4}16^{-3/4}=\frac{1}{4}2^{-3}=\frac{1}{4}\frac{1}{8}=\frac{1}{32}\] \[L(x)=f(16)+f'(16)(x-16)=2+\frac{1}{32}(x-16)\] \[L(16.04)=2+\frac{1}{32}(16.04-16)=2+\frac{1}{32}\cdot .04=2.00125\]